# Gradient descent induces alignment between weights and the empirical NTK for deep non-linear networks

### v1

@@ Appendix C Extending our theoretical predictions to depth and general activations
Figure 8: Observed versus the first-order predicted C-NFA for the input to the first layer of a two hidden layer MLP. The dashed line is neural network width $k=n=d=128$, where $n$ and $d$ are the number of data point and data dimension, respectively, while the solid line uses $n=d=128$ and $k=1024$.
Appendix D Proofs and derivations
Proof of Proposition 5.
Consider the second derivatives of the NFM and the AGOP.
$\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}\bar{W})$ $=\dot{W}^{\top}\bar{W}+\bar{W}^{\top}\dot{W}~{}.$
Then,
$\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})$ $=2\dot{W}^{\top}\dot{W}+\ddot{W}^{\top}\bar{W}+\bar{W}^{\top}\ddot{W}~{}.$
At initialization $\bar{W}=0$, therefore,
$\bar{W}^{\top}\bar{W}=0,\quad\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}\bar{W})=0,\quad\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})=2\dot{W}^{\top}\dot{W}~{}.$ (6)
Meanwhile, for the (centered) AGOP,
$\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}K\bar{W})$ $=\dot{W}^{\top}K\bar{W}+\bar{W}^{\top}\dot{K}_{z}\bar{W}+\bar{W}^{\top}K\dot{W}~{}.$
Then,
$\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}K\bar{W})$ $=\ddot{W}^{\top}K\bar{W}+\dot{W}^{\top}\dot{K}\bar{W}+\dot{W}^{\top}K_{z}\dot{W}$
$+\dot{W}^{\top}K\dot{W}+\bar{W}^{\top}\dot{K}\dot{W}+\bar{W}^{\top}K\ddot{W}$
$+\dot{W}^{\top}\dot{K}\bar{W}+\bar{W}^{\top}\ddot{K}\bar{W}+\bar{W}^{\top}\dot{K}\dot{W}~{}.$
At initialization,
$\bar{W}^{\top}K\bar{W}=0,$ $\quad\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}K\bar{W})=0,$ (7)
$\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}K\bar{W})$ $=2\dot{W}^{\top}K\dot{W}~{}.$ (8)
Therefore, the correlation between the centered neural feature matrix and the centered AGOP at initialization is determined by,
$\rho\left(\dot{W}^{\top}\dot{W},\dot{W}^{\top}K\dot{W}\right)~{}.$
Both sides simplify at initialization:
$\dot{W}^{\top}\dot{W}=X^{\top}YKYX,\quad\dot{W}^{\top}K\dot{W}=X^{\top}YK^{2}YX~{},$
where $K$ is its value at initialization. ∎
Appendix E Free probability calculations of C-NFA
In order to understand the development of the NFA, we analyze the centered NFA in the limit that learning rate is much smaller than the initialization for a one hidden layer MLP with quadratic activations. We write this particular network as,
@@ Appendix E Free probability calculations of C-NFA
where $a\in\mathbb{R}^{1\times k}$ and $W\in\mathbb{R}^{k\times d}$, where $d$ is the input dimension and $k$ is the width. In this case, the NFA has the following form,
@@
We assume that computing the average of the NFA quantities across initializations is equal to the quantities themselves in the high-dimensional limit.
@@
We will compute the expected values of the centered NFA under these assumptions. In the remainder of the section we will drop the $\mathbb{E}\left[\cdot\right]$ in the trace for ease of notation.
@@ E.2 Numerator term of NFA
E.3 First denominator term of NFA
@@
E.4 Second denominator term of NFA
@@
All terms of the NFA are now in terms of traces of the matrices $A$, $B$, and $R$ and functions on each term separately. The matrices $A$ and $B$ are determined by the data, while the moments of the eigenvalues of $R$ are determined by the initialization distribution of the weights in the neural network, and neither training nor the data.
Appendix F Balance dataset
The data consists of a mixture of two distributions from which two subsets of the data $X_{1}$ and $X_{2}$ are sampled from, and is parametrized by a balance parameter $\gamma\in(0,1]$ and two variance parameters $\epsilon_{1},\epsilon_{2}>0$. The subset $X_{1}$ which has label $y_{1}=1$ and constitutes a $\gamma$ fraction of the entire dataset, is sampled from a multivariate Gaussian distribution with mean 00 and covariance
@@ Appendix F Balance dataset
Then the second subset, $X_{2}$, is constructed such that $X_{2}^{\top}X_{2}\approx(X_{1}^{\top}X_{1})^{-2}$, and has labels $y_{2}=0$. Then, for balance parameter $\gamma$ sufficiently small, the AGOP second derivative approximately satisfies,
$=X_{1}^{\top}X_{1}X^{\top}XX_{1}^{\top}X_{1}$
$\approx X_{1}^{\top}X_{1}X_{2}^{\top}X_{2}X_{1}^{\top}X_{1}$
$\approx I~{},$
In contrast, the NFM second derivative, $\mathbb{E}\left[\dot{W}^{\top}\dot{W}\right]=(X^{\top}YX)^{2}=(X_{1}^{\top}X_{1})^{2}\approx\Sigma^{2}$, will be significantly far from identity.
Motivated by this derivation, we construct $X_{2}$ by the following procedure:
@@
    Extract singular values $S_{1}$ and right singular vectors $U_{1}$ from a singular-value decomposition (SVD) of $X_{1}^{\top}X_{1}$.
@@
    Extract the left singular vectors $V_{2}$ from a sample $\tilde{X}_{2}$ that is sampled from the same distribution as $X_{1}$.
@@
    Where $X=X_{1}\oplus X_{2}$, Set $X\leftarrow X+\epsilon_{2}Z$, where $Z\sim\mathcal{N}(0,I)$.
@@
Note that $U_{1}S_{1}^{-1}V_{2}^{\top}V_{2}S_{1}^{-1}U_{1}^{\top}=U_{1}S_{1}^{-2}U_{1}^{\top}=(X_{1}^{\top}X_{1})^{-2}$, therefore, we should set $X_{2}=V_{2}S_{1}^{-1}U_{1}^{\top}$ to get $X_{2}^{\top}X_{2}=(X_{1}^{\top}X_{1})^{-2}$. Regarding the variance parameters, in practice we set $\epsilon_{1}=0.5$ and $\epsilon_{2}=10^{-2}$.
@@
[…]
@@ Appendix G Varying the data distribution
We verify that our observations for isotropic Gaussian data hold even when the data covariance has a significant spectral decay. (Figures 9 and 10). We again consider Gaussian data that is mean 00 and where the covariance is constructed from a random eigenbasis. In Figure 9, we substitute the eignevalue decay as $\lambda_{k}\sim\frac{1}{1+k}$, while in Figure 10, we use $\lambda_{k}\sim\frac{1}{1+k^{2}}$. We plot the values of the UC-NFA, C-NFA, train loss, and test loss throughout training for the first and second layer of a two hidden layer network with ReLU activations, while additionally varying initialization scale. Similar to Figure 2, we observe that the C-NFA is more robust to the initialization scale than the UC-NFA, and UC-NFA value become high through training, while being small at initialization. We see that the test loss improves for smaller initializations, where the value of the C-NFA and UC-NFA are higher.
Figure 9: Centered NFA correlations. Data covariance decay rate $\lambda_{k}\sim\frac{1}{1+k}$. Top row is layer 1, bottom row is layer 2. Train (test) losses are scaled by the maximum train (test) loss achieved so that they are between 00 and $1$.
Figure 10: Centered NFA correlations. Data covariance decay rate $\lambda_{k}\sim\frac{1}{1+k^{2}}$. Top row is layer 1, bottom row is layer 2. Train (test) losses are scaled by the maximum train (test) loss achieved so that they are between 00 and $1$.
@@ Appendix H Effect of initialization on feature learning
We see that when initialization is small, the C-NFA and UC-NFA are high at the end of training with and without fixing the learning speed (Figure 4). This is reflected by the quality of the features learned by the NFM and the qualitative similarity of the NFM and AGOP at small initialization scale (Figure 11). Further, we notice that as we increase initialization, without fixing speeds, the correspondence between the NFM and decreases and the quality of the NFM features decreases (at a faster rate than the AGOP). Strikingly, when learning speeds are fixed, the quality of the features in the AGOP and NFM becomes invariant to the initialization scale.
@@ Appendix I Experimental details
For the experiments in Figure 1, we used $n=384$ data points, $d=32$, $k=128$ as the width in all layers, isotropic Gaussian data, initialization scale $0.01$ in the first layer and default scale in the second. We used ReLU activations and two hidden layers. For the experiments in Figure 2,9,10,6, and 7, we used a two hidden layer network with ReLU activations, learning rate $0.05$, 800 steps of gradient decent, and took correlation/covariance measurements every 5 steps.
@@
For the SLO figures (Figures 5, 11, 4), we use isotropic Gaussian data, 600 steps of gradient descent. The learning rates are chosen based on initialization scale in the first layer. For initialization scales 1, 0.1, 0.01, and 0.001, we used learning rates $0.03$, $0.1$, $0.2$, $0.4$, respectively. We again used two hidden layers with ReLU activations. We chose $n=256$, $d=32$, and $k=256$ as the width. We divided the linear readout weights by $0.01$ at initialization to promote feature learning, and modified SLO to scale gradients by $(\epsilon+\|\nabla\mathcal{L}\|)^{-1}$, rather than just the inverse of the norm of the gradient, for $\epsilon=0.1$. This technique smooths the training dynamics as the parameters approach a loss minimum, allowing the network to interpolate the labels.
@@
For the Deep C-NFA predictions (Figure 8), we used $n=128$, $d=128$, initialization scale of $1$. The low rank task is just the chain monomial of rank $r=5$. The high rank polynomial task is $y(x)=\sum_{i=1}^{d}(Qx)_{i}^{2}$, where $Q\in\mathbb{R}^{d\times d}$ is a matrix with standard normal entries.
Real datasets
For the experiments on the SVHN dataset, we train a four hidden layer neural network with ReLU activations, initialization scale $1.0$ in all layers and width $256$. For SVHN, we subset the dataset to 4000 points. We train for 3000 epochs with learning rate 0.2 for standard training, and 0.3 for SLO, and take NFA measurements every 50 epochs. For SLO, we set $C_{0}=2.5$, $C_{1}=C_{2}=0.4$, and relaxation parameter $\epsilon=0.2$. We pre-process the dataset so that each pixel is mean 00 and standard deviation $1$. For the experiments on CelebA, we train a two hidden layer network on a balanced subset of 7500 points with Adam with learning rate $0.0001$ and no weight decay. We use initialization scale 0.02 in the first layer, and width 128. We train for 500 epochs. We pre-process the dataset by scaling the pixel values to be between 0 and 1.
@@
We demonstrate the SLO can be applied adaptively to increase the strength of the UC-NFA in all layers of a deep network on the chain monomial task of rank $r=3$. We train a three hidden layer MLP with ReLU activations and an initialization scale of $0.1$ by SLO, and find that all layers finish at the same high UC-NFA (Figure 12). Further, this final UC-NFA value is higher than the highest UC-NFA achieved by any layer with standard training. The generalization loss is also lower with SLO on this example, corresponding to better feature learning (through the UC-NFA).
Figure 12: Training with SLO where the learning speeds are chosen adaptively based on the UC-NFA values of all layers. The dashed lines correspond to training with standard GD.
At every time step we choose $C_{i}=s$ for the layer $i$ with the smallest UC-NFA correlation value, while setting $C_{j}=s^{-1}$ for all other layers, with $s=20$. We again modify SLO by dividing the gradients by ϵitalic-$\epsilon\|\nabla\mathcal{L}\|$ for $\epsilon=0.01$. The learning rate is set to $0.05$ in SLO and $0.25$ for the standard training (gradient descent), and the networks are trained for 500 epochs. We sample $n=256$ points with $d=32$, and use width $k=256$.
Appendix K Experiments on real datasets
We replicate Figures 1, 4, and 5 on celebrity faces (CelebA) and Street View House Numbers (SVHN). We begin by showing that one can disrupt the NFA correspondence by replacing the PTK feature covariance with a random matrix of the same spectral decay. For this example, we measure the Pearson correlation, which subtracts the mean of the image. I.e. $\bar{\rho}(A,B)\equiv\rho\left(A-m(A),B-m(B)\right)$, where $m(A),m(B)$ are the average of the elements of $A$ and $B$.
@@
Figure 14: We demonstrate on the SVHN dataset, with a 4 hidden layer neural network with large initialization scale, how SLO can improve the strength of the UC-NFA, the C-NFA, the ratio of the unnormalized C-NFA to UC-NFA (plot (a)) and the feature quality (plots (b) and (c)). In plots (b) and (c), we visualize the diagonal of the NFM and AGOP for the first layer of the trained network, where SLO was applied with $C_{0}=2.5$, $C_{1}=C_{2}=0.4$.

### v2

@@ Appendix D Extending our theoretical predictions to depth and general activations
Figure 7: Observed versus the first-order predicted C-NFC for the input to the first layer of a two hidden layer MLP. The dashed line is neural network width $k=n=d=128$, where $n$ and $d$ are the number of data point and data dimension, respectively, while the solid line uses $n=d=128$ and $k=1024$.
Appendix E Free probability calculations of C-NFC
In order to understand the development of the NFC, we analyze the centered NFC in the limit that learning rate is much smaller than the initialization for a one hidden layer MLP with quadratic activations. We write this particular network as,
@@ Appendix E Free probability calculations of C-NFC
where $a\in\mathbb{R}^{1\times k}$ and $W\in\mathbb{R}^{k\times d}$, where $d$ is the input dimension and $k$ is the width. In this case, the NFC has the following form,
@@
We assume that computing the average of the NFC quantities across initializations is equal to the quantities themselves in the high-dimensional limit.
@@
We will compute the expected values of the centered NFC under these assumptions. In the remainder of the section we will drop the $\mathbb{E}\left[\cdot\right]$ in the trace for ease of notation.
@@ E.2 Numerator term of NFC
E.3 First denominator term of NFC
@@
E.4 Second denominator term of NFC
@@
All terms of the NFC are now in terms of traces of the matrices $A$, $B$, and $R$ and functions on each term separately. The matrices $A$ and $B$ are determined by the data, while the moments of the eigenvalues of $R$ are determined by the initialization distribution of the weights in the neural network, and neither training nor the data.
Appendix F Alignment reversing dataset
The data consists of a mixture of two distributions from which two subsets of the data $X^{(1)}$ and $X^{(2)}$ are sampled from, and is parametrized by a balance parameter $\gamma\in(0,1]$ and two variance parameters $\epsilon_{1},\epsilon_{2}>0$. The subset $X^{(1)}$ which has label $y_{1}=1$ and constitutes a $\gamma$ fraction of the entire dataset, is sampled from a multivariate Gaussian distribution with mean 00 and covariance
@@ Appendix F Alignment reversing dataset
We construct $X^{(2)}$ by the following procedure:
@@
    Extract singular values $S_{1}$ and right singular vectors $U_{1}$ from a singular-value decomposition (SVD) of $(X^{(1)})^{\top}X^{(1)}$.
@@
    Extract the left singular vectors $V_{2}$ from a sample $\tilde{X}_{2}$ that is sampled from the same distribution as $X^{(1)}$.
@@
    Where $X=X^{(1)}\oplus X^{(2)}$, Set $X\leftarrow X+\epsilon_{2}Z$, where $Z\sim\mathcal{N}(0,I)$.
@@
Note that $U_{1}S_{1}^{-1}V_{2}^{\top}V_{2}S_{1}^{-1}U_{1}^{\top}=U_{1}S_{1}^{-2}U_{1}^{\top}=((X^{(1)})^{\top}X^{(1)})^{-2}$, therefore, we should set $X^{(2)}=V_{2}S_{1}^{-1}U_{1}^{\top}$ to get $(X^{(2)})^{\top}X^{(2)}=((X^{(1)})^{\top}X^{(1)})^{-2}$. Regarding the variance parameters, in practice we set $\epsilon_{1}=0.5$ and $\epsilon_{2}=10^{-2}$.
@@
$\mathbb{E}_{a,W}\left[\dot{W}^{\top}K\dot{W}\right]$ $=3\cdot\text{tr}\left(X^{\top}X\right)\cdot(X^{\top}YX)^{2}$
@@ Appendix G Varying the data distribution
We verify that our observations for isotropic Gaussian data hold even when the data covariance has a significant spectral decay. (Figures 9 and 10). We again consider Gaussian data that is mean 00 and where the covariance is constructed from a random eigenbasis. In Figure 9, we substitute the eignevalue decay as $\lambda_{k}\sim\frac{1}{1+k}$, while in Figure 10, we use $\lambda_{k}\sim\frac{1}{1+k^{2}}$. We plot the values of the UC-NFC, C-NFC, train loss, and test loss throughout training for the first and second layer of a two hidden layer network with ReLU activations, while additionally varying initialization scale. Similar to Figure 8, we observe that the C-NFC is more robust to the initialization scale than the UC-NFC, and UC-NFC value become high through training, while being small at initialization. We see that the test loss improves for smaller initializations, where the value of the C-NFC and UC-NFC are higher.
Figure 8: Centered neural feature correlations. Data covariance decay rate $\lambda_{k}=1$. Top row is layer 1, bottom row is layer 2. Train (test) losses are scaled by the maximum train (test) loss achieved so that they are between 00 and $1$.
Figure 9: Centered neural feature correlations. Data covariance decay rate $\lambda_{k}\sim\frac{1}{1+k}$. Top row is layer 1, bottom row is layer 2. Train (test) losses are scaled by the maximum train (test) loss achieved so that they are between 00 and $1$.
Figure 10: Centered neural feature correlations. Data covariance decay rate $\lambda_{k}\sim\frac{1}{1+k^{2}}$. Top row is layer 1, bottom row is layer 2. Train (test) losses are scaled by the maximum train (test) loss achieved so that they are between 00 and $1$.
@@ Appendix H Effect of initialization on feature learning
We see that when initialization is small, the C-NFC and UC-NFC are high at the end of training with and without fixing the learning speed (Figure 4). This is reflected by the quality of the features learned by the NFM and the qualitative similarity of the NFM and AGOP at small initialization scale (Figure 11). Further, we notice that as we increase initialization, without fixing speeds, the correspondence between the NFM and decreases and the quality of the NFM features decreases (at a faster rate than the AGOP). Strikingly, when learning speeds are fixed, the quality of the features in the AGOP and NFM becomes invariant to the initialization scale.
@@ Appendix I Experimental details
For the experiments in Figure 1, we used $n=384$ data points, $d=32$, $k=128$ as the width in all layers, isotropic Gaussian data, initialization scale $0.01$ in the first layer and default scale in the second. We used ReLU activations and two hidden layers. For the experiments in Figure 8,9,10,5, and 6, we used a two hidden layer network with ReLU activations, learning rate $0.05$, 800 steps of gradient decent, and took correlation/covariance measurements every 5 steps.
C/UC-NFC calculations on real datasets
We describe the experimental details for Figure 2. For (A,B) we trained a five layer MLP on the first 50,000 datapoints of Streetview House Numbers (SVHN), CIFAR-10, CIFAR-100, STL-10, MNIST, and German Traffic Sign Recognition Benchmark (GTSRB) datasets. We used the default PyTorch initialization (scale of $1$) for all layers. We used width 256 in all layers, and trained with SGD with batch size 128. For SVHN and CIFAR, we trained for 150 epochs with learning rate 0.2. For STL-10 and GTSRB we used learning rate 0.1 for 150 epochs. For MNIST we trained for 50 epochs with learning rate 10. For the VGG-11 experiments on CIFAR-10, we used the default architecture from torchvision with batch-norm layers removed. We trained for 50 epochs and learning rate 1. For experiments with a GPT-family model, we adapted the model and dataset from NanoGPT (https://github.com/karpathy/nanoGPT). We used all the default settings and the default Adam optimizer, but used no weight decay, learning rate 5e-3, and removed all dropout layers. We also reduced the number of attention layers to 3 from the original 6. We trained on the Shakespeare characters dataset.
@@
For the SLO figures (Figures 11, 4), we use isotropic Gaussian data, 600 steps of gradient descent. The learning rates are chosen based on initialization scale in the first layer. For initialization scales 1, 0.1, 0.01, and 0.001, we used learning rates $0.03$, $0.1$, $0.2$, $0.4$, respectively. We again used two hidden layers with ReLU activations. We chose $n=256$, $d=32$, and $k=256$ as the width. We divided the linear readout weights by $0.01$ at initialization to promote feature learning, and modified SLO to scale gradients by $(\epsilon+\|\nabla\mathcal{L}\|)^{-1}$, rather than just the inverse of the norm of the gradient, for $\epsilon=0.1$. This technique smooths the training dynamics as the parameters approach a loss minimum, allowing the network to interpolate the labels.
@@
For the Deep C-NFC predictions (Figure 7), we used $n=128$, $d=128$, initialization scale of $1$. The low rank task is just the chain monomial of rank $r=5$. The high rank polynomial task is $y(x)=\sum_{i=1}^{d}(Qx)_{i}^{2}$, where $Q\in\mathbb{R}^{d\times d}$ is a matrix with standard normal entries.
Figures 16 and 17
For the experiments on the SVHN dataset, we train a four hidden layer neural network with ReLU activations, initialization scale $1.0$ in all layers and width $256$. For SVHN, we subset the dataset to 4000 points. We train for 3000 epochs with learning rate 0.2 for standard training, and 0.3 for SLO, and take NFC measurements every 50 epochs. For SLO, we set $C_{0}=2.5$, $C_{1}=C_{2}=0.4$, and relaxation parameter $\epsilon=0.2$. We pre-process the dataset so that each pixel is mean 00 and standard deviation $1$. For the experiments on CelebA, we train a two hidden layer network on a balanced subset of 7500 points with Adam with learning rate $0.0001$ and no weight decay. We use initialization scale 0.02 in the first layer, and width 128. We train for 500 epochs. We pre-process the dataset by scaling the pixel values to be between 0 and 1.
Code availability
We make the code available for the experiments in Figure 2 available at https://anonymous.4open.science/r/centered_NFA-4795/ (for MLP+VGG).
@@
We demonstrate the SLO can be applied adaptively to increase the strength of the UC-NFC in all layers of a deep network on the chain monomial task of rank $r=3$. We train a three hidden layer MLP with ReLU activations and an initialization scale of $0.1$ by SLO, and find that all layers finish at the same high UC-NFC (Figure 12). Further, this final UC-NFC value is higher than the highest UC-NFC achieved by any layer with standard training. The generalization loss is also lower with SLO on this example, corresponding to better feature learning (through the UC-NFC).
Figure 12: Training with SLO where the learning speeds are chosen adaptively based on the UC-NFC values of all layers. The dashed lines correspond to training with standard GD.
At every time step we choose $C_{i}=s$ for the layer $i$ with the smallest UC-NFC correlation value, while setting $C_{j}=s^{-1}$ for all other layers, with $s=20$. We again modify SLO by dividing the gradients by ϵitalic-$\epsilon\|\nabla\mathcal{L}\|$ for $\epsilon=0.01$. The learning rate is set to $0.05$ in SLO and $0.25$ for the standard training (gradient descent), and the networks are trained for 500 epochs. We sample $n=256$ points with $d=32$, and use width $k=256$.
Appendix K Additional C-NFC/UC-NFC plots across architectures
Figure 13: Full trajectories for the C/UC-NFC of a five hidden layer MLP trained on six datasets, averaged over three seeds, with large ($1.0$) initialization scale. The blue curves are the UC-NFC, while green curves are the C-NFC.
Figure 14: Losses and normalized changes in weights across datasets for a five hidden layer MLP. The change in weight is measured as $\|W-W_{0}\|\|W\|^{-1}$. First column of all subfigures are the losses, while the second columns are the weight changes.
Figure 15: Losses and normalized changes in weights across datasets for VGG-11 on CIFAR-10 and GPT on Shakespeare character generation. The change in weight is measured as $\|W-W_{0}\|\|W\|^{-1}$.
First column of (A) and (B) are the losses, while the remaining columns are the weight changes.
Appendix L Additional experiments on real datasets
We replicate Figures 1 and 4 on celebrity faces (CelebA) and Street View House Numbers (SVHN). We begin by showing that one can disrupt the NFC correspondence by replacing the PTK feature covariance with a random matrix of the same spectral decay. For this example, we measure the Pearson correlation, which subtracts the mean of the image. I.e. $\bar{\rho}(A,B)\equiv\rho\left(A-m(A),B-m(B)\right)$, where $m(A),m(B)$ are the average of the elements of $A$ and $B$.
@@
Figure 17: We demonstrate on the SVHN dataset, with a 4 hidden layer neural network with large initialization scale, how SLO can improve the strength of the UC-NFC, the C-NFC, the ratio of the unnormalized C-NFC to UC-NFC (plot (a)) and the feature quality (plots (b) and (c)). In plots (b) and (c), we visualize the diagonal of the NFM and AGOP for the first layer of the trained network, where SLO was applied with $C_{0}=2.5$, $C_{1}=C_{2}=0.4$.